# One flow to correct them all: improving simulations in high-energy physics with a single normalising flow and a switch

### v1

@@ 1 Introduction
An alternative is to address the underlying mismodelling of the relevant observables, such as the features that are used to reconstruct a certain particle. Such an approach promises to be more general, as any algorithm that is based on the corrected features is expected to show a better agreement of simulation and data. One approach is to derive weights for simulated events. These can be obtained from a classifier that is trained to distinguish simulation and data [1, 2, 3]. However, a disadvantage of the weighting approach is that observables that are not part of the training may show a worse agreement between simulations and data after the weights are applied. The weighting approach may also lead to increased uncertainties due to the limited number of samples in the simulation in case of large weights. Another approach is to morph the observables in the simulation to correspond to the multidimensional distribution observed in data so that only the targeted observables are modified. This approach has been implemented with chained quantile regressions [4], generative adversarial networks [5], input convex neural networks [6] for optimal transport properties [7], and normalising flows [8, 9, 10]. In the latter approach, normalising flows [11, 12] are trained as a bijective transformation between the complex multidimensional distribution in the input space and a simple “base distribution” of the same dimensionality, often a multivariate Gaussian distribution. Since monotonically increasing bijective transformations are used, normalising flows preserve the quantiles of the distributions, which makes them suitable for morphing.
@@ 2.1 Normalising flows for morphing distributions
Normalising flows are a class of generative models designed for the effective learning and sampling of multivariate probability distributions. They are constructed by parametrising a transformation that performs the mapping of a complex input distribution to a tractable base distribution of the same dimension, $d$. Such a transformation, $f$, must be invertible and hence ensure a one-to-one correspondence between the input probability density function (PDF), $p_{x}(\mathbb{x})$, and the base PDF, $p_{z}(\mathbb{z})$, i.e. $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$. Given that the composition of multiple invertible functions remains invertible, it is common practice to construct the transformation $f$ as a composition of $N$ invertible transformations to increase the expressiveness of the model [12, 13], ${f=f_{1}\circ\ldots\circ f_{N}}$.
@@
We used masked autoregressive flows [14] based on spline transformations [15, 16, 17, 18]. The autoregressive property of the flow leads to an efficient computation of the Jacobian determinant. The auxiliary networks are implemented as multilayer perceptrons with masked connections, for which we use MADE blocks [19] (Masked Autoencoder for Distribution Estimation). Using splines in normalising flows has the advantage that they can approximate more complex distributions compared to affine transformations while still being simple to invert. We use neural spline flows [18], which are based on monotonic rational quadratic splines. The monotonicity of the transformation is important for morphing distributions, because they ensure that the quantiles of the distributions are preserved during the transformation.
@@ 2.2 Two-dimensional benchmarks
We use the neural spline implementation of the Zuko package [21] in PyTorch [22]. We set the number of splines to $8$. For the MADE blocks, we use two hidden layers with $128$ nodes each and ReLU activation. We use the ADAM optimiser [23] with a cosine annealing learning-rate scheduler [24]. The training is performed until the validation loss does not improve over $15$ epochs. For the two-dimensional benchmark examples, the normalising flow is composed of four monotonic rational-quadratic transformations.
Figure 2 shows the result of the morphing, where each plot displays the original distributions on the left with the resulting morphed distributions on the right. Our approach is able to reproduce also the sharp edges and discontinuous features in both sets of distributions. We note that it is simple to extend the boolean condition to a one-hot encoding in order to morph between more than two domains with a single flow. We illustrate this in a three-domain example in Appendix B.
@@ 3 Generation of the physics-inspired dataset
As a final step, the seven variables are endowed with non-trivial correlations using the mcerp package [25, 26]. While the marginal distributions before and after the artificial correlation are very similar, we are able to impose non-trivial correlations between all features, which increases the complexity of the dataset. The resulting correlation matrices differ significantly between data and simulation. Figure 3 shows the marginal distributions of the seven variables for data and simulation.
We generate two statistically independent datasets for toy data and toy simulation, respectively, with ten million events each. More technical details for the generation of the dataset are given in Appendix A. The code for generating the dataset is publicly available [^1] https://github.com/JaLuka98/OneFlowPublic.
Figure 3: Marginal distributions for the seven variables in the data and simulation datasets. The four informative features are shown in the upper figures and the three ancillary variables in the lower figures. The ancillary variables are defined in a way that $p_{\mathrm{T}}$ is unitless, $\eta$ takes only positive values and $N$ is defined in the interval [0,3].
@@ 4.1 Preprocessing and training
Figure 4 illustrates the forward pass of the normalising flow for the example of the informative feature $v^{\mathrm{B}}_{1}$. In the forward pass, all four informative features are transformed to the four-dimensional base distribution via autoregressive rational quadratic splines, as described in Section 2. The MADE blocks are conditioned on the IsData boolean and the three ancillary variables $p_{\mathrm{T}}$, $\eta$ and $N$.
@@
Discontinuous informative features are subject to a smoothing step, making them continuous. The reason is that normalising flows use differentiable transformations to map the original distribution to a differentiable base distribution. As this can only be approximate for discontinuous input distributions, the overall performance of normalising flows for the morphing application may suffer in such cases. In our dataset, this applies to the variables $v^{\mathrm{B}}_{1}$ and $v^{\mathrm{B}}_{2}$.
@@ 4.2 Evaluation of the corrections
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 5 Conclusions
The morphing approach with a single normalising flow shows excellent performance and seems a promising tool for the correction of complex simulated distributions in high-energy physics and beyond. Finally, we note that it is simple to extend the boolean switch to a multiclass condition to morph between more than two domains (cf. Appendix B).
@@ Acknowledgements
Appendix A Details of the dataset generation
@@
Appendix B Morphing between three domains

### v2

@@ Abstract
keywords:
Domain adaptation, Normalising flows, High energy physics, Monte Carlo corrections
@@ 1 Introduction
An alternative is to address the underlying mismodelling of the relevant observables, such as the features that are used to reconstruct a certain particle. Such an approach promises to be more general, as any algorithm that is based on the corrected features is expected to show a better agreement of simulation and data. One approach is to derive weights for simulated events. These can be obtained from a classifier that is trained to distinguish simulation and data [1, 2, 3, 4]. However, a disadvantage of the weighting approach is that observables that are not part of the training may show a worse agreement between simulations and data after the weights are applied. The weighting approach may also lead to increased uncertainties due to the limited number of samples in the simulation in case of large weights. Another approach is to morph the observables to correspond to the multidimensional target distribution so that only certain observables are modified. This approach has been implemented with chained quantile regressions [5], generative adversarial networks [6], input convex neural networks [7] for optimal transport properties [8], generative diffusion networks [9] and normalising flows [10, 11, 12]. In the latter approach, normalising flows [13, 14] are trained as a bijective transformation between the complex multidimensional distribution in the input space and a simple “base distribution” of the same dimensionality, often a multivariate Gaussian distribution. Since monotonically increasing bijective transformations are used, normalising flows preserve the quantiles of the distributions, which makes them suitable for morphing.
@@ 2.1 Normalising flows for morphing distributions
Normalising flows are a class of generative models designed for the effective learning and sampling of multivariate probability distributions. They are constructed by parametrising a transformation that performs the mapping of a complex input distribution to a tractable base distribution of the same dimension, $d$. Such a transformation, $f$, must be invertible and hence ensure a one-to-one correspondence between the input probability density function (PDF), $p_{x}(\mathbf{x})$, and the base PDF, $p_{z}(\mathbf{z})$, i.e. $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$. Given that the composition of multiple invertible functions remains invertible, it is common practice to construct the transformation $f$ as a composition of $N$ invertible transformations to increase the expressiveness of the model [14, 15], ${f=f_{1}\circ\ldots\circ f_{N}}$.
@@
We used masked autoregressive flows [16] based on spline transformations [17, 18, 19, 20]. The autoregressive property of the flow leads to an efficient computation of the Jacobian determinant. The auxiliary networks are implemented as multilayer perceptrons with masked connections, for which we use MADE blocks [21] (Masked Autoencoder for Distribution Estimation). Using splines in normalising flows has the advantage that they can approximate more complex distributions compared to affine transformations while still being simple to invert. We use neural spline flows [20], which are based on monotonic rational quadratic splines. The monotonicity of the transformation is important for morphing distributions, because they ensure that the quantiles of the distributions are preserved during the transformation. When a transformation is monotonic, the order of data points is maintained. This means that the same proportion of data points will fall below any given threshold in both the original and transformed distributions. Consequently, the quantiles, which represent these thresholds, remain unchanged.
@@ 2.2 Two-dimensional benchmarks
We use the neural spline implementation of the Zuko package [23] in PyTorch [24]. We set the number of bins in each spline to $8$. For the MADE blocks, we use two hidden layers with $128$ nodes each and ReLU activation. We use the ADAM optimiser [25] with a cosine annealing learning-rate scheduler [26]. The training is performed until the validation loss does not improve over $15$ epochs. For the two-dimensional benchmark examples, the normalising flow is composed of four monotonic rational-quadratic transformations.
Figure 2 shows the result of the morphing, where each plot displays the original distributions on the left with the resulting morphed distributions on the right. Our approach is able to reproduce also the sharp edges and discontinuous features in both sets of distributions. We note that it is simple to extend the boolean condition to a one-hot encoding in order to morph between more than two domains with a single flow. We illustrate this in a three-domain example in Appendix C.
@@ 3 Generation of the physics-inspired dataset
As a final step, the seven variables are endowed with non-trivial correlations using the mcerp package [27, 28]. While the marginal distributions before and after the artificial correlation are very similar, we are able to impose non-trivial correlations between all features, which increases the complexity of the dataset. The resulting correlation matrices differ significantly between data and simulation. Figure 3 shows the marginal distributions of the seven variables for data and simulation. Two-dimensional visualisations with $68\%$ contours are given in Appendix A.
We generate two statistically independent datasets for toy data and toy simulation, respectively, with ten million events each. More technical details for the generation of the dataset are given in Appendix B. The code for generating the dataset is publicly available [^1] https://github.com/JaLuka98/OneFlowPublic.
Figure 3: Marginal distributions for the seven variables in the data and simulation datasets. The three ancillary variables are shown in the upper figures and the four informative features in the lower figures. The ancillary variables are defined in a way that $p_{\mathrm{T}}$ is unitless, $\eta$ takes only positive values and $N$ is defined in the interval [0,3].
@@ 4.1 Preprocessing and training
Figure 4 illustrates the forward pass of the normalising flow for the example of the informative feature $v^{\mathrm{B}}_{1}$. In the forward pass, all four informative features are transformed to the four-dimensional base distribution via autoregressive rational quadratic splines, as described in Section 2. The MADE blocks are conditioned on the IsData boolean and the three ancillary variables $p_{\mathrm{T}}$, $\eta$ and $N$. Thus, the learning objective is the conditional probability $p(v^{\mathrm{A}}_{1},v^{\mathrm{B}}_{1},v^{\mathrm{A}}_{2},v^{\mathrm{B}}_{2}\mid p_{\mathrm{T}},\eta,N,\mathrm{isData})$. This setup corresponds to the typical application of this correction method in high-energy physics, where the model is trained in a control region and applied in a signal region with different distributions of the ancillary features.
@@
Discontinuous informative features are subject to a smoothing step, making them continuous. The reason is that normalising flows use differentiable transformations to map the original distribution to a differentiable base distribution. As this can only be approximate for discontinuous input distributions, the overall performance of normalising flows for the morphing application may suffer in such cases [29]. In our dataset, this applies to the variables $v^{\mathrm{B}}_{1}$ and $v^{\mathrm{B}}_{2}$.
@@ 5 Conclusions
The morphing approach with a single normalising flow shows excellent performance and seems a promising tool for the correction of complex simulated distributions in high-energy physics and beyond. Finally, we note that it is simple to extend the boolean switch to a multiclass condition to morph between more than two domains (cf. Appendix C).
@@ Acknowledgements
Appendix A Two-dimensional visualisation of the dataset
Figure 12 visualises the physics-inspired dataset by showing two-dimensional 68% contours of each pair of variables. This shows their non-trivial correlations and highlights that these are different between toy simulation and toy data.
Figure 12: Visualisation of the physics-inspired dataset in selected ranges for simulation (blue) and data (orange). On the diagonal, the marginal distributions are shown. On the off-diagonal, the $68\%$ contours obtained from kernel density esimation are depicted. The ancillary variables are defined in a way that $p_{\mathrm{T}}$ is unitless, $\eta$ takes only positive values and $N$ is defined in the interval [0,3].
Appendix B Details of the dataset generation
@@
Appendix C Morphing between three domains